OpenStudy (anonymous):

3 sin theta=sin theta-1

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

$$ 3sin(\theta)=sin(\theta)-1 \implies 2sin(\theta)=-1 $$ what would you get for \(sin(\theta)\) there?

OpenStudy (anonymous):

im not sure how to get that?

OpenStudy (jdoe0001):

ok, lemme rewrite it 3a = a -1 3a (-a) = a(-a )-1 2a = -1 what would "a" be?

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

what if I were to put it this way 2x = -1 what would "x" be?

OpenStudy (anonymous):

okay lets get this straight that thing called a theta er whatever its after sin...what is that to be exact?

OpenStudy (jdoe0001):

is an angle

OpenStudy (jdoe0001):

can you get the "x" or "a" in the above equations? once you get that, you pretty much get the angle :)

OpenStudy (anonymous):

hope this helps